# Collet–Eckmann maps in the unicritical family

### v1

@@ 4 Promotion
We see that
$\frac{\tilde{\alpha}_{\nu_{j}}(c)}{\underline{\gamma}_{\nu_{j}}(c)}\leq\kappa^{\prime j}\frac{\tilde{\alpha}_{\nu_{0}}(A)}{\underline{\gamma}_{\nu_{0}}(A)},\qquad\text{ where $c\in A$}.$
@@
where $C$ is a distortion constant (close to $1$). Since $\kappa^{\prime}>\kappa$, we have $\kappa^{\prime s+1}-1-\kappa+\kappa^{-s}\geq\kappa^{\prime}-\kappa>0$ for all $s\geq 0$. This finishes the proof since $\alpha_{\nu}\nu=r$ is large (so that the constant $4C$ in front is eaten up). ∎
@@
So the remaining parameters constitute a finite union of squares, were the $\alpha=\alpha_{\nu_{J}}(c)$ are small, sufficiently small in order to use the large deviation argument. Note that $\delta_{1}$, the portion deleted, can be made arbitrarily small. Heuristically, the method of promotion indicates that all CE-maps are density points of slowly recurrent maps, since one can choose $\alpha>0$ arbitrarily small (possibly diminishing the diameter of ${\mathcal{Q}}$).
5 Conclusion
We are going to consider iterates from some starting time $m_{0}$ (not necessarily $N$) up to $(1+\iota)m_{0}$ for some suitable $\iota>0$. As in previous papers, [7, 8] and [4] we define escape time as follows.
Definition 5.1.
5.1 Large deviations
In this section we choose $\tilde{C}$ so that $\alpha_{\nu_{J}}/\underline{\gamma}_{\nu_{J}}$ is small enough. Let us see how the Lyapunov exponent changes during the time of promotion. During the promotion period between $N=\nu_{0}$ and $\nu_{J}=m_{0}$ the Lyapunov exponent cannot drop too much, since all the bound periods $p_{j}$ are bounded by $p_{0}$, by Lemma 3.6. But we have something stronger. Recall that $\underline{\gamma}_{j}$ is also defined for $j<N$, namely $\underline{\gamma}_{j}=\underline{\gamma}$, and $\gamma_{p_{j}}\geq(d-1)\alpha_{p_{j}+1}+\underline{\gamma}_{p_{j}}$, where $\underline{\gamma}_{p_{j}}=\underline{\gamma}$ since $p_{j}\leq p_{0}<N$. By Lemma 3.3, we have, if $10<\eta<$ 1,
@@ 5.1 Large deviations
In the following lemma, we estimate the maximum escape time for essential returns after promotion in an interval $[m_{0},2m_{0}]$. Note that during this period, the exponent cannot drop below $\gamma_{I}/4$.
@@
where $M=2d/(\gamma_{C}(1-\hat{\kappa}))$. Moreover, $p_{0}\leq 8dr/\gamma_{I}$ (since the exponent $\gamma\geq\gamma_{I}/4)$, so
@@
By Lemma 3.9, we can assume without loss of generality that $\xi_{N}({\mathcal{Q}})$ is an essential return. Otherwise, if $\xi_{N}({\mathcal{Q}})$ is in escape position, we can go directly to the next section. The point is that most parameters in ${\mathcal{Q}}$ escape before a certain time after $N$. Suppose $m_{0}=\nu_{J}$ is an essential return time for a partition element $A=A_{0}$. We now want to choose $\iota$ so that the Lyapunov exponent cannot drop under $\gamma_{C}$ anytime in the time window $[m_{0},(1+\iota)m_{0}]$.
@@
where $\eta(\Delta)\rightarrow 0$ as $\Delta\rightarrow\infty$. Let $A_{s,R}$ be the set of parameters in $A_{0}$ that have exactly $s$ free essential returns in the time interval $[m_{0},(1+\iota)m_{0}]$ before escape, where $R=r_{1}+\ldots+r_{s}$. Let $\hat{A}$ be the largest of all squares in $A_{s,R}$. Then,
@@
This means that a big portion of $A_{0}$ will reach the large scale within the time period $[m_{0},(1+\iota)m_{0}]$, where $\iota=10\tilde{M}\frac{3+\hat{\kappa}}{1-\hat{\kappa}}\alpha\leq 1$ (see the condition on $\alpha$ in (5.3)). Let this portion be $1-\delta_{2}$, where $10<\delta_{2}<$ 1, is small. Again note that $\delta_{2}$ can be chosen as close to zero as we want.
@@
5.2 Hyperbolic maps
Now to prove our theorems, we can argue similarly as in [4]. Let ${\mathcal{F}}(f_{c_{0}})$ be the Fatou set for the unperturbed map $f_{c_{0}}$ which is Collet–Eckmann, and let ${\mathcal{N}}_{\varepsilon_{0}}$ be an $\varepsilon_{0}$-neighbourhood of ${\mathcal{J}}(f_{c_{0}})$ as defined in (2.8). Here, $\varepsilon_{0}$ can be chosen arbitrarily close to zero since the Hausdorff dimension of the Julia set of $f_{c_{0}}$ is strictly less than two and therefore has zero area [11, Corollary 2]. Then since the Fatou set of $f_{c_{0}}$ consists only of the attracting basin of $\infty$, any compact subset of the Fatou set is stable under perturbation. So ^$\widehat{{\mathbb{C}}}\setminus{\mathcal{N}}_{\varepsilon_{0}}$ is still contained the Fatou set ${\mathcal{F}}(f_{c})$ for all $c\in{\mathcal{Q}}$, provided that ${\mathcal{Q}}$ is chosen small enough. By bounded distortion
@@ 5.2 Hyperbolic maps
where $\delta_{3}$ is a constant depending on the distortion of $\xi_{n}(c)$ on ${\mathcal{Q}}$. Since also $10<\delta_{3}<$ 1 can be arbitrarily small, it follows that the Lebesgue density of hyperbolic maps at $c=c_{0}$ is equal to $1$, by making the diameter of ${\mathcal{Q}}$ tending to zero.
@@
The methods of promotion and the following parameter exclusion are likely to generalise to all rational CE-maps. Since the recurrence exponent $\alpha>0$ after the promotion period can be chosen arbitrarily small (with possibly diminishing the diameter of ${\mathcal{Q}}$), taking into consideration that there may be several critical points in the Julia set (see [3], [4]), one could prove that every rational CE-map for which $J(f)=\hat{{\mathbb{C}}}$ is a Lebesgue density point of slowly recurrent CE-maps, even for real-analytic families of maps containing the starting CE-map (cf. [17] where complex one-dimensional families were considered).

### v2

@@ 4 Promotion
We see that, if $A$ is a partition element at time $\nu_{j-1}$,
$\frac{\alpha_{\nu_{j}}(c)}{\gamma_{\nu_{j}}(c)}\leq\frac{\tilde{\alpha}_{\nu_{j}}(A)}{\underline{\gamma}_{\nu_{j}}(A)}\leq\kappa^{\prime j}\frac{\tilde{\alpha}_{\nu_{0}}(A)}{\underline{\gamma}_{\nu_{0}}(A)},\qquad\text{ where $c\in A$}.$
@@
where $C$ is a distortion constant (close to $1$). Since $\kappa^{\prime s+1}+\kappa^{-s}$ increases in $s\geq 0$ we have $\kappa^{\prime s+1}-1-\kappa+\kappa^{-s}\geq\kappa^{\prime}-\kappa>0$ for all $s\geq 0$. This finishes the proof since $\alpha_{\nu}\nu\sim r$ is large (so that the constant $4C$ in front is eaten up). ∎
@@
So the remaining parameters constitute a finite union of squares, were the $\alpha=\alpha_{\nu_{J}}(c)$ are sufficiently small in order to use the large deviation argument. Note that $\delta_{1}$, the portion deleted, can be made arbitrarily small by decreasing the size of $\mathcal{Q}$}. Heuristically, the method of promotion indicates that all CE-maps are density points of slowly recurrent maps, since one can choose $\alpha>0$ arbitrarily small (possibly diminishing the diameter of $\mathcal{Q}$}).
5 After promotion
We are going to consider iterates from some starting time $m_{0}=\nu_{J}$, which is the end of the promotion period (it could be $N$, but not necessarily) up to $(1+\iota)m_{0}$ for some suitable $\iota>0$.
5.1 Uniform bound of the promotion period
In this section we choose $\tilde{C}$ so that $\alpha_{\nu_{J}}/\underline{\gamma}_{\nu_{J}}$ is small enough. Let us see how the Lyapunov exponent changes during the time of promotion. During the promotion period between $N=\nu_{0}$ and $\nu_{J}=m_{0}$ the Lyapunov exponent cannot drop too much, since all the bound periods $p_{j}$ are bounded by $p_{0}$, by Lemma 3.6. But we have something stronger. Recall that $\underline{\gamma}_{j}$ is also defined for $j<N$, namely $\underline{\gamma}_{j}=\underline{\gamma}$, and $\gamma_{p_{j}}\geq(d-1)\alpha_{p_{j}+1}+\underline{\gamma}_{p_{j}}$, where $\underline{\gamma}_{p_{j}}=\underline{\gamma}$ since $p_{j}\leq p_{0}<N$. By Lemma 3.3, we have, if $0<\eta<1$,
@@ 5.1 Uniform bound of the promotion period
5.2 Large deviations
As in previous papers, [9, 10] and [5] we define escape time as follows.
Definition 5.2.
In the following lemma, we estimate the maximum escape time for essential returns after promotion in an interval $m_{0},2m_{0}]$. Note that during this period, the exponent cannot drop below $\gamma_{I}/4$.
@@ 5.2 Large deviations
where $)M=2d/(\gamma_{C}(1-\hat{\kappa}$)). Moreover, $p_{0}\leq 8dr/\gamma_{I}$ (since the exponent $\gamma\geq\gamma_{I}/4)$, so
@@
By Lemma 3.9, we can assume without loss of generality that $\xi_{N}({\mathcal{Q}})$ is an essential return. Otherwise, if $\xi_{N}({\mathcal{Q}})$ is in escape position, we can go directly to the next section. The point is that most parameters in $\mathcal{Q}$} escape before a certain time after $N$. Suppose $m_{0}=\nu_{J}$ is an essential return time for a partition element $A=A_{0}$. We now want to choose $\iota$ so that the Lyapunov exponent cannot drop under $\gamma_{C}$ anytime in the time window $m_{0},(1+\iota)m_{0}]$.
@@
where $\eta(\Delta)\rightarrow 0$ as $\Delta\rightarrow\infty$. Let $A_{s,R}$ be the set of parameters in $A_{0}$ that have exactly $s$ free essential returns in the time interval $m_{0},(1+\iota)m_{0}]$ before escape, where $R=r_{1}+\ldots+r_{s}$. Let $\hat{A}$ be the largest of all squares in $A_{s,R}$. Then,
@@
This means that a big portion of $A_{0}$ will reach the large scale within the time period $m_{0},(1+\iota)m_{0}]$, where $\iota=10\tilde{M}\frac{3+\hat{\kappa}}{1-\hat{\kappa}}\alpha\leq 1$ (see the condition on $\alpha$ in (5.3)). Let this portion be $1-\delta_{2}$, where $0<\delta_{2}<1$, is small. Again note that $\delta_{2}$ can be chosen as close to zero as we want.
@@
6 Hyperbolic maps
Now to prove our theorems, we can argue similarly as in [5]. Let ${\mathcal{F}}(f_{c_{0}})$ be the Fatou set for the unperturbed map $f_{c_{0}}$ which is Collet–Eckmann, and let ${\mathcal{N}}_{\varepsilon_{0}}$ be an $\varepsilon_{0}$-neighbourhood of ${\mathcal{J}}(f_{c_{0}})$ as defined in (2.8). Here, $\varepsilon_{0}$ can be chosen arbitrarily close to zero since the Hausdorff dimension of the Julia set of $f_{c_{0}}$ is strictly less than two and therefore has zero area [15, Corollary 2]. Then since the Fatou set of $f_{c_{0}}$ consists only of the attracting basin of $\infty$, any compact subset of the Fatou set is stable under perturbation. So $\widehat{{\mathbb{C}}}\setminus{\mathcal{N}}_{\varepsilon_{0}}$ is still contained the Fatou set ${\mathcal{F}}(f_{c})$ for all $c\in{\mathcal{Q}}$, provided that $\mathcal{Q}$} is chosen small enough. By bounded distortion
@@ 6 Hyperbolic maps
where $\delta_{3}$ is a constant depending on the distortion of $\xi_{n}(c)$ on $\mathcal{Q}$}. Since also $0<\delta_{3}<1$ can be arbitrarily small, it follows that the Lebesgue density of hyperbolic maps at $c=c_{0}$ is equal to $1$, by making the diameter of $\mathcal{Q}$} tending to zero.
@@
The methods of promotion and the following parameter exclusion are likely to generalise to all rational CE-maps. Since the recurrence exponent $\alpha>0$ after the promotion period can be chosen arbitrarily small (with possibly diminishing the diameter of $\mathcal{Q}$}), taking into consideration that there may be several critical points in the Julia set (see [4][5]), one could prove that every rational CE-map for which $J(f)=\hat{{\mathbb{C}}}$ is a Lebesgue density point of slowly recurrent CE-maps, even for real-analytic families of maps containing the starting CE-map (cf. [23] where complex one-dimensional families were considered).